OpenStudy (anonymous):

which expression is equivalent to (1-sin^2 x)(sec x)? sec^2 x tan x tan^2 x cos x cos^2 x

OpenStudy (anonymous):

i don't know

OpenStudy (butterflydreamer):

hint: \[1 = \sin^2x + \cos^2x , secx = \frac{ 1 }{ cosx }\]

OpenStudy (anonymous):

what is this.what grade are you on

OpenStudy (anonymous):

11th grade algebra 2

OpenStudy (anonymous):

for real

OpenStudy (anonymous):

I'm going to 9 grade next year

OpenStudy (butterflydreamer):